# Observing Hot Holographic Quark Star With Gravitational Waves

### v1

@@ 4 Conclusions
Applying the five connected EoS, we calculate the M-R relation and the I-Love-Q-C relations. Our analysis on the M-R relation, along with Table.(1) which lists the maximum mass-radius values for both the core and the whole of each model, reveal that most models have heavy cores, leading to maximum masses that are similar to those of "simple” QS. The I-Love-Q-C panels show violations for "simple” QS especially with $\mu/T=4.5$. However, the presence of thin hadron shells sharing the same EoS mitigates these violations, as the outer layers of realistic NS EoS exhibit similarities.
The results have significant implications for future observations. Firstly, the maximum masses of the hot QS are around 23 to 30 $\rm{M_{\odot}}$, which suggest that they could be observed in GW events as BH mimickers. Since BH exhibit zero TLN in GR, one could distinguish hot QS from BH through the inference on TLN. Moreover, the deviation from universality in the I-Love-Q-C relations for "simple” QS provides a means to infer the underlying quark model. Comparing with the I-Love-Q-C relations for the "simple” QS, the universal relations for the hybrid models remain valid, enabling the potential discrimination between them, verifying the hypothesis in Wu:2023aaz that the outer layer contributes most to the universal relations. The universal relations could also be employed to constrain the parameters in the phase transition.
@@
For the future work, one quick direction is to consider the 2-flavor QCD case discussed in Zhao:2023gur . On the other hand, we will pay more attention on the detailed structure of the hot hadron outer layer. Moreover, it will be interesting to study the constructions of these hybrid stars when changing the outer layer EoS.
Acknowledgements. The authors thank Song He, Chian-Shu Chen, Alessandro Parisi, Chen Zhang and Zhoujian Cao for very helpful discussions. KZ (Hong Zhang) is supported by a classified fund from Shanghai city.

### v2

@@ 4 Conclusions
Applying the five connected EoS, we calculate the M-R relation and the I-Love-Q-C relations. Our analysis on the M-R relation, along with Table.(1) which lists the maximum mass-radius values for both the core and the whole of each model, reveal that most models have heavy cores, leading to maximum masses that are similar to those of "simple" QS. The I-Love-Q-C panels show violations for "simple" QS especially with $\mu/T=4.5$. However, the presence of thin hadron shells sharing the same EoS mitigates these violations, as the outer layers of realistic NS EoS exhibit similarities.
The results have significant implications for future observations. First, the maximum masses of the hot QS are around 23 to 30 $\rm{M_{\odot}}$, which suggest that they could be observed in GW events as BH mimickers. Since BH exhibit zero TLN in GR, one could distinguish hot QS from BH through the inference on TLN. Moreover, the deviation from universality in the I-Love-Q-C relations for "simple" QS provides a means to infer the underlying quark model. Comparing with the I-Love-Q-C relations for the "simple" QS, the universal relations for the hybrid models remain valid, enabling the potential discrimination between them, verifying the hypothesis in Wu:2023aaz that the outer layer contributes most to the universal relations. The universal relations could also be employed to constrain the parameters in the phase transition.
@@
For the next step, one quick direction is to consider the 2-flavor QCD case discussed in Zhao:2023gur , with some progress given in Chen:2025ppr . On the other hand, we will pay more attention on the detailed structure of the hot hadron outer layer. Moreover, it will be interesting to study the constructions of these hybrid stars when changing the outer layer EoS.
The formation mechanism of hot QS and their lifetime are also interesting. One possibility is that it originates from the high-temperature early universe, while other possibilities include the collapse of a massive star, along with the binary mergers. Since a supernova explosion or binary merger can reach a temperature of 50 MeV Sekiguchi:2011mc ; Drago:2015dea , there is a good chance to hold a hot QS if this magnitude could double, achieving a quark state. This state could resist gravity through thermal motion and degeneracy pressure, preventing the formation of a BH. In fact, its relatively large radius results in a compactness of approximately $0.1$, which is lower than the typical NS compactness of $0.2$ and much smaller than the BH compactness of $0.5$.
Considering the high temperature of the quark core, the radiation seems to be quite strong. However, the model we are discussing is not a bare quark star but rather one where the quark state forms the core, while the shell consist of hadronic matter that has undergone a phase transition. As we move toward the star’s surface, the temperature gradually decreases, leading to a reduction in the star’s overall radiation. On the other hand, at the boundary between the core and the outer layers (i.e., the phase transition region), radiation can be reflected back, maintaining a high-temperature core and a low-temperature outer shell. Radiation might indeed cause such a QS to have a relatively short lifespan, which could explain why no observational evidence has been found so far (or they might have been mistaken for BH, and due to their distance, tidal deformations or electromagnetic effects have not been detected). As the temperature falls, there will be a second collapse, which can be distinguished from a standard one-time explosion supernova. We can further discuss the possible existence time of this type of QS. The details are beyond the scope of the current paper, and will be left for future work.
Acknowledgements. The authors thank Song He, Chian-Shu Chen, Alessandro Parisi, Chen Zhang and Zhoujian Cao for very helpful discussions. K.Z. (Hong Zhang) is supported by a classified fund from Shanghai city.